A box contains 3 pens 2 markers and 1 highlighter.sara selects one item

at random and does not return it to the box she then selects a second

item at random what is the probability that sara selects one pen and

then one marker.


1
Expert's answer
2021-06-28T18:42:26-0400

Given,

3 pens, 2 markers and 1 highlighters


Total = 6


Probability that sara selects one pen and

then one marker-


"P=\dfrac{^3C_1\times ^2C_1}{^6C_2}"


"=\dfrac{6}{15}=0.4"